OpenStudy (anonymous):

there are 785 students in a senior class. there are 77 more females than males. how many male and female seniors are there?

OpenStudy (anonymous):

f + m = 785 m = 77 + f

OpenStudy (anonymous):

subsstitue and solve with the above equations

OpenStudy (anonymous):

77+ f + f = 785 77 + 2f = 785 2f = 708 f = 354

OpenStudy (anonymous):

m = 77 + 354 m = 431

OpenStudy (anonymous):

check 431 + 354 = 785 :)

OpenStudy (anonymous):

f=m+77 f+m= m+77+m = 2m + 77 2m + 77 =785 please tell me u can do basic algebra

OpenStudy (anonymous):

juanitaM: m does not equal f+77

OpenStudy (anonymous):

785 students in a senior class. there are 77 more females f + m = 785 f - m = 77 -> 2f = 862 => f = 431 females => m = 354 males